OpenStudy (anonymous):

(PLEASE HELP) Three consecutive odd integers are such that the square of the third integer I s9 less than the sum of the squares of the first two. One solution is -3, -1, and 1. Find three other consecutive odd integers that also satisfy the given conditions. ( I keep getting the wrong answer)

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello.

OpenStudy (misty1212):

i assume you solved \[n^2+(n+2)^2=(n+4)^2+9\]

OpenStudy (anonymous):

Honestly I don't know since I don't understand what exactly is wrong

OpenStudy (misty1212):

ok lets go from the beginning first number is \(n\) then the second would be \(n+2\) and the third would be \(n+4\) right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

so sum of squares of the first two are \[n^2+(n+2)^2\] and 9 more than the third is \((n+4)^2+9\)

OpenStudy (misty1212):

you get this quadratic equation \[n^2+(n+2)^2=(n+4)^2+9\] which will have two solutions

OpenStudy (misty1212):

one is \(-3\) so \(-3,-1,1\) is one answer

OpenStudy (misty1212):

the other is \(7\) so \(7,9,11\) we can solve it if you like

OpenStudy (anonymous):

yes please, so -3 and 7 are two of the answers?

OpenStudy (misty1212):

yeah \[n^2+(n+2)^2=(n+4)^2+9\] expand get \[n^2+n^2+4n+4=n^2+8n+16\]

OpenStudy (misty1212):

oops forgot the 9 !

OpenStudy (misty1212):

\[n^2+n^2+4n+4=n^2+8n+16+9\] thats better

OpenStudy (misty1212):

then put all on one side get \[n^2-4n-21=0\] factor and you get your two solutions

OpenStudy (anonymous):

so there would be ONLY two solutions right?

OpenStudy (misty1212):

two solutions for the first number \(n\)

OpenStudy (anonymous):

(x-3)(x+7) would turn to 3,-7?

OpenStudy (misty1212):

yeah it would but to factor correctly it is \[(x-7)(x+3)=0\]

OpenStudy (anonymous):

ohh alright I got it from here thank you very much.